. Assume that arrivals occur according to a Poisson process with an average


of seven per hour. What is the probability that exactly two customers arrive in the two-hour period


of time between


(a) 2:00 P.M. and 4:00 P.M. (one continuous two-hour period)?


(b) 1:00 P.M. and 2:00 P.M. or between 3:00 P.M. and 4:00 P.M. (two separate one-hour periods


that total two hours)?



1
Expert's answer
2022-03-29T01:08:25-0400

We have a Poisson distribution,

"\lambda=7;\\
P_t(X=k)=\cfrac{(\lambda t)^k\cdot e^{-\lambda t}}{k!}."


"(a) \ \ P_{2}(X=2)=\cfrac{(7\cdot2)^2\cdot e^{-7\cdot2}}{2!}=\\
=0.00008."


"(b) \ \ P_{1}(X=2)\cdot P_{1}(X=2)=\\
=\begin{pmatrix}
 \cfrac{(7\cdot1)^2\cdot e^{-7\cdot1}}{2!}
\end{pmatrix}^2=\\
=0.02234^2=0.00050."